a 71kg diver drops from a board 10m above the water surface. use conservation of mechanical energy to find his speed 5m above the water surface.


1
Expert's answer
2021-11-10T10:10:22-0500

"E_0=P+K,"

"E_0=mgH,"

"P=mgh,"

"K=\frac{mv^2}2,\implies"

"v=\sqrt{2g(H-h)}=10~\frac ms."